OpenStudy (anonymous):

Not sure on how to solve this problem. The function t(x) = 3x + 5 determines how many cans of corn kernels a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on corn. The function c(x) = 2x + 3. Solve for how much money must be spent when the crew is going to work 3 shifts. Answers: 11 14 31 43 Is it A or B?

jimthompson5910 (jim_thompson5910):

x = 3 shifts t(x) = 3x + 5 t(3) = 3*3 + 5 ... replace x with 3 t(3) = ??? then plug the result you get for t(3) into c(x)

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome